# LANTERN: Accelerating Visual Autoregressive Models with Relaxed Speculative Decoding

### v1

@@ 4.2 Main Results
Table 3: Actual speed-up, mean accepted length, FID, and CLIP score for each method. $\tau$ refers to the temperature for a generation. The actual speed-up is measured on a single A100 80GB SXM.
@@
Speedup Mean Accepted Length FID CLIP score
Vanilla AR (Sun et al., 2024) 1.00$\times$ 1.00 28.63 0.3169
EAGLE-2 (Li et al., 2024a) 1.28$\times$ 1.60 - -
LANTERN ($\delta=0.05$, $k=1000$) 1.68$\times$ 2.02 29.77 0.3164
LANTERN ($\delta=0.20$, $k=1000$) 2.25$\times$ 2.89 30.78 0.3154
@@
Speedup Mean Accepted Length FID CLIP score
Vanilla AR (Sun et al., 2024) 1.00$\times$ 1.00 15.22 0.3203
EAGLE-2 (Li et al., 2024a) 0.93$\times$ 1.20 - -
LANTERN ($\delta=0.10$, $k=1000$) 1.17$\times$ 1.75 16.17 0.3208
LANTERN ($\delta=0.40$, $k=1000$) 1.65$\times$ 2.40 18.76 0.3206
Figure 4: Qualitative samples for LANTERN and vanilla autoregressive decoding. From top to bottom, the images are generated by vanilla, LANTERN ($\delta=0.05$, $\delta=0.1$, $\delta=0.2$, $\delta=0.4$) where $k$ is fixed at 1000, and images in the same column are generated using the same text prompt.
@@ 4.2.1 Speed Up Comparison
To confirm that LANTERN provides notable speed improvements over the baseline while maintaining image quality, we compare our method with baselines under both greedy decoding and random sampling situations. Table 3 demonstrates the speedup and image quality across different methods. LANTERN shows a significant acceleration, even with some degradation in image quality, when compared to vanilla autoregressive decoding and EAGLE-2 (Li et al., 2024a).
For $\tau=0$, LANTERN achieves a 2.25×2.25$\times$2.25 × actual speedup with a slight quality cost, as reflected in a 2.16 increase in FID. In contrast, EAGLE-2 only achieves 1.28×1.28$\times$1.28 × speedup, which is far less than ours. LANTERN demonstrates that by slightly relaxing the (i.e. $\delta=0.2$), we can achieve higher speedup compared to other baselines while maintaining the acceptable quality loss, with an FID of 30.78, which is close to the vanilla AR decoding.
For $\tau=1$, while EAGLE-2 fails to achieve actual speedup, our approach accomplishes 1.65×1.65$\times$1.65 × speedup by relaxing the acceptance condition, with a cost of a 3.53 increase in FID. Nevertheless, a more conservative setting of $\delta=0.1$ yields a balanced compromise, which obtains 1.17×1.17$\times$1.17 × speedup with only a 6% increase in FID compared to the vanilla AR, making it a suitable choice when quality preservation is critical. CLIP score is maintained within a similar level, regardless of $k$ and $\delta$, which indicates that LANTERN does not alter alignment between input text and generated images.
Similar to the other speculative decoding methods in LLMs, we can achieve greater speedup with greedy decoding compared to the sampling. However, vision autoregressive models often fail to generate high quality images with greedy decoding, which leads to the inferior FID score. As a result, by allowing a degree of flexibility in token selection, our method strikes a favorable balance between speed and quality, outperforming both vanilla AR and EAGLE-2 in terms of practical efficiency.
@@ 4.2.2 Qualitative Results
To confirm that image quality is preserved with LANTERN, as indicated by the FID score, we conduct qualitative analysis. Figure 4 demonstrates that, despite the modification of the target model’s probability distribution to achieve acceleration, our method effectively preserves image quality. Notably, even under the setting of $\delta=0.4$ and $k=1000$, which achieve about 1.6×1.6$\times$1.6 × speedup compared to the vanilla decoding, generated images retain both content and style at a level comparable to standard autoregressive decoding. These qualitative results, along with the fact that LANTERN avoids significant degradation in FID score as shown in the previous section, demonstrate that it effectively preserves image quality while increasing efficiency. More qualitative examples can be found in Appendix D.
@@ 4.2.3 Trade-off Between Performance and Efficiency
LANTERN provides various options between quality and efficiency to the end users. Therefore, we explore this trade-off by adjusting $k$ and $\delta$ across different settings. Figure 5 illustrates the relationship between image quality, which is measured by FID, and speedup, which is assessed with mean accepted length for various settings of $\delta$ and $k$, under both sampling ($\tau=1$) and greedy decoding ($\tau=0$). The trade-off curves highlight that increasing $\delta$ and $k$ generally improves speedup, but at the expense of image quality.
In the case of greedy decoding, we observe that while larger $\delta$ tends to increase the speedup, it may not always result in substantial acceleration when $k$ is small. This could be attributed to the inherent nature of greedy decoding, which only accepts the token with the top-1 probability. If $k$ is not large enough, the increase in acceptance probability is insufficient to surpass the top-1 threshold, thus limiting the speedup gains.
For sampling, note that we observe different behaviors. When $\delta$ is small such as 0.05 and 0.1, increasing $k$ maintains image quality while still achieving better acceleration. This suggests that by tuning $\delta$ and selecting an appropriate $k$, we can find an optimal trade-off where the speedup is improved without significantly compromising image quality. Specifically, larger $k$ values allow for faster generation without disproportionately degrading FID, particularly at lower $\delta$ values.
These findings confirm that LANTERN allows flexible tuning of performance versus efficiency, where careful adjustment of $\delta$ and $k$ can yield substantial acceleration with minimal quality loss. By setting appropriate hyperparameters, particularly under sampling, we can achieve significant speedups while preserving the quality of images, offering a better balance compared to the original speculative decoding method.
@@ 4.3 Ablation Study
In this section, we conduct ablation studies for LANTERN. In Section 4.3.1, we assess the impact of the metric used to measure latent proximity on performance. Then, in Section 4.3.2, we provide an ablation study on the effect of the metric for measuring distance from the modified probability distribution.
@@ 4.3.1 Nearest Latent Selection
Table 4: Ablation study for latent proximity measure and probability distribution distance metrics. Best performing method annotated with bold, and second best is marked with underline.
@@
TVD ($\delta=0.4$) 2.40 18.76 0.3206
To explore the impact of various metrics for measuring latent proximity, we conduct an ablation study using representative distance metrics commonly used to measure latent proximity. Table 4 summarizes the comparisons among different strategies for selecting the nearest latent tokens, including $\ell_{2}$ distance, cosine similarity, and random selection. This experiment aims to assess the role of proximity-based selection in token aggregation, with $k=1000$ used across all methods.
As shown in Table 4, the random selection significantly underperforms in terms of acceleration, achieving only a $1.26$ mean accepted length, which is notably lower than both $\ell_{2}$ distance and cosine similarity. Additionally, the random selection shows inferior acceleration compared to the $\ell_{2}$ distance with $\delta=0.05$, while maintaining a similar FID, which highlights the importance of token selection based on latent proximity.
Comparing $\ell_{2}$ distance and cosine similarity, both methods demonstrate comparable performance, suggesting robustness in our approach to proximity measurement. Mean accepted length is almost identical between these two, and FID/CLIP score also have a marginal gap. This suggests that our method is effective regardless of the specific proximity metric considered, offering flexibility without significant trade-offs in performance. These results confirm that selecting tokens based on latent proximity plays a crucial role in both acceleration and image quality, with proximity-based metrics like $\ell_{2}$ distance and cosine similarity, providing a clear advantage over random selection.
@@ 4.3.2 Distance Between Probability Distribution
To ensure that the modified target distribution remains within an acceptable range of divergence from the original, we introduce $\delta$ as an upper bound for distributional divergence. To validate the impact of the divergence metric, we evaluate two different metrics to measure this divergence: Total Variation Distance (TVD) and Jensen-Shannon Divergence (JSD). Kullback-Leibler Divergence (KLD) is not used as it is asymmetric and not a valid metric for a distance in mathematical sense.
To compare the effectiveness of these distance metrics, we fix $k$ as 1000 and adjust $\delta$ to achieve similar mean accepted lengths across the different methods. The results, shown in Table 4, demonstrate that JSD with $\delta=0.2$ results in a performance that achieves between TVD with $\delta=0.2$ and with $\delta=0.4$. Specifically, JSD yields 2.29 mean accepted length with 18.21 FID, positioning it between two TVD configurations in terms of both speed and image quality.
These results confirm that our method consistently functions as a robust trade-off controller regardless of the chosen distance metric. Since the difference in performance between two metrics is marginal, we opt to use TVD, as it is computationally lighter and thus more efficient for large-scale implementations.
@@ Appendix A Related Works
Visual Autoregressive Models
Based on the advancements of autoregressive (AR) learning in natural language processing (Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023), autoregressive models have been extended to generative vision tasks by tokenizing images into discrete tokens arranged on a 2D grid and defining a unidirectional token sequence. Early studies (Esser et al., 2021; Yu et al., 2021; Lee et al., 2022) employ row-major raster scan, spiral, or z-curve orders to generate sequences, but these approaches are computationally inefficient and demonstrated inferior performance compared to diffusion models. In response, recent visual AR models (Sun et al., 2024; Tian et al., 2024) introduce a novel tokenizer and multi-scale, coarse-to-fine ordering strategy, establishing a scalable paradigm that surpasses diffusion models. More recently, Chameleon (Team, 2024) propose a multi-modal AR model capable of generating both images and text within a unified framework, demonstrating strong versatility across diverse tasks.
Speculative Decoding for LLMs
LLM inference is known to be memory-bounded, which means that the computation (or processing) is blocked by the slow transfer of data (e.g., model parameters) (Leviathan et al., 2023). To address the issue, the speculative decoding (Leviathan et al., 2023) has been proposed to make more computations at a single decoding step by attaching the draft sequences after the usual input tokens. Speculative decoding theoretically guarantee that prediction results with speculative decoding exactly match with target models distribution. Speculative decoding uses a smaller model which is trained on the same dataset as a drafter, thus it involves significant extra latency by running another model (drafter) and it hinders the actual acceleration.
Medusa (Cai et al., 2024) proposed a light-weighted drafter where the size is similar to the lm_head of the target model and suggested to use somewhat lenient acceptance condition called typical acceptance rather than speculative sampling. Although the typical acceptance does not provide any fundamental guarantee about the resulting distribution, they show that the results’ quality is maintained compared to the target model. EAGLE (Li et al., 2024a) proposes to make the drafter slightly heavier than Medusa by utilizing single autoregressive layers, but it dramatically improves the acceleration.
In this work, we mainly follow the EAGLE as a base method since it exhibits the best performance in LLM literature. Based on its its drafter architecture and training settings, we trained drafter on image data and adapt it for speculative decoding.
Multi-modal speculative decoding
Recently, Gagrani et al. (2024) proposes a speculative decoding method for the multi-modal large language models (MLLMs). They mainly focus on the model, which processes both images and texts as input and generates texts such as LLaVA. Note that we focus on the visual autoregressive models regarding image generation in this work, which is a totally different model with the scope of Gagrani et al. (2024).
Appendix B Algorithm

### v3

@@ 5.2 Main Results
Table 2: Actual speed-up, MAL (Mean Accepted Length), FID, CLIP score, Precision / Recall, and HPS v2 for each method. $\tau=0$ refers to the greedy decoding and $\tau=1$ refers to the sampling with temperature $1$ for generation. The actual speed-up is measured on a single RTX 3090.
@@
Acceleration ($\uparrow$) Image Quality Metrics
Speed-up MAL FID ($\downarrow$) CLIP score ($\uparrow$) Precision ($\uparrow$) Recall ($\uparrow$) HPSv2 ($\uparrow$)
Vanilla AR (Sun et al., 2024) 1.00$\times$ 1.00 28.63 0.3169 0.4232 0.3517 23.18
EAGLE-2 (Li et al., 2024a) 1.29$\times$ 1.60 - - - - -
LANTERN ($\delta=0.05$, $k=1000$) 1.56$\times$ 2.02 29.77 0.3164 0.4484 0.3158 22.62
LANTERN ($\delta=0.2$, $k=1000$) 2.26$\times$ 2.89 30.78 0.3154 0.4771 0.2773 21.69
@@
Acceleration ($\uparrow$) Image Quality Metrics
Speed-up MAL FID ($\downarrow$) CLIP score ($\uparrow$) Precision ($\uparrow$) Recall ($\uparrow$) HPSv2 ($\uparrow$)
Vanilla AR (Sun et al., 2024) 1.00$\times$ 1.00 15.22 0.3203 0.4781 0.5633 24.11
EAGLE-2 (Li et al., 2024a) 0.93$\times$ 1.20 - - - - -
LANTERN ($\delta=0.1$, $k=1000$) 1.13$\times$ 1.75 16.17 0.3208 0.4869 0.5172 23.75
LANTERN ($\delta=0.4$, $k=1000$) 1.69$\times$ 2.40 18.76 0.3206 0.4909 0.4497 23.22
Figure 4: Qualitative samples generated by LlamaGen-XL Stage II model for LANTERN and standard autoregressive decoding. From top to bottom, the images are generated by standard autoregressive decoding, LANTERN ($\delta=0.2$, $\delta=0.4$) where $k$ is fixed at 1000, and images in the same column are generated using the same text prompt. Text prompts for the images are provided in Appendix I.
@@ 5.2.1 Speed Up Comparison
To confirm that LANTERN provides notable speed improvements over the baseline while maintaining image quality, we compare our method with baselines under both greedy decoding and random sampling situations. Table 2 demonstrates the speed-up and image quality across different methods. LANTERN shows a significant acceleration, even with a slight degradation in image quality, when compared to standard autoregressive decoding and EAGLE-2 (Li et al., 2024a). Our method, LANTERN demonstrates significant improvements in both mean accepted length and actual speed-up. In terms of mean accepted length, LANTERN outperforms both baselines by reaching 2.40×2.40$\times$2.40 × and it translate to substantial actual speed-up in practice, with LANTERN achieving 1.69×1.69$\times$1.69 × actual speed-up.
While LANTERN’s acceleration comes with a trade-off in image quality, the degradation remains minimal and well within acceptable boundaries. For instance, the HPS v2 score decreases by less than $1.5$ for both greedy decoding and sampling when compared to standard autoregressive decoding. Precision remains stable or slightly improved, whereas recall shows a modest decline, indicating a slight reduction in image diversity but with individual image quality largely preserved. Other metrics, such as FID and CLIP score, further confirm that LANTERN maintains competitive image quality despite the significant acceleration.
As a result, by allowing a degree of flexibility in token selection, our method strikes a favorable balance between speed and quality, outperforming both standard autoregressive decoding and EAGLE-2 in terms of practical efficiency. Experimental results on other visual AR models including LlamaGen-XL Stage II model and Anole (Chern et al., 2024) are presented in Appendix E. Also, detailed analyses of latency, including latency on Intel Gaudi 2, are provided in Appendix F.
@@ 5.2.2 Qualitative Results
To confirm that image quality is preserved with LANTERN, as indicated by the various image quality metrics, we conduct a qualitative analysis. Figure 4 demonstrates that, despite the modification of the target model’s probability distribution to achieve acceleration, our method effectively preserves image quality. Notably, even under the setting of $\delta=0.4$ and $k=1000$, which achieve about 1.64×1.64$\times$1.64 × speed-up compared to the standard autoregressive decoding, generated images retain both content and style at a level comparable to standard autoregressive decoding. These qualitative results, along with the fact that LANTERN avoids significant degradation in image quality metrics as shown in the previous section, demonstrate that it effectively preserves image quality while increasing efficiency. More qualitative examples can be found in Appendix J.
@@ 5.2.3 Trade-off Between Performance and Efficiency
LANTERN provides various options between quality and efficiency to the end users. Therefore, we explore this trade-off across different hyperparameters by adjusting $k$ and $\delta$. Figure 5 illustrates the relationship between image quality, which is measured by FID, and speed-up, which is assessed with mean accepted length for various settings of $\delta$ and $k$, under both sampling ($\tau=1$) and greedy decoding ($\tau=0$). The trade-off curves highlight that increasing $\delta$ and $k$ generally improves speed-up, but at the expense of image quality.
The results underline that LANTERN allows for flexible tuning to balance acceleration and image quality depending on specific practical requirements. For instance, in the sampling ($\tau=1$), configurations with smaller $\delta$ or $k$ prioritize image quality, maintaining lower FID scores, while larger $\delta$ or $k$ achieve greater acceleration with a controlled loss in quality. A similar trend is observed for greedy decoding ($\tau=0$), where the trade-off is more pronounced at higher mean accepted lengths. Importantly, across all hyperparameter configurations, LANTERN demonstrates consistent and predictable trade-offs, enabling users to fine-tune the method based on their preferred balance between speed and quality. Trade-offs with other image quality metrics can be found in Appendix H.
@@ 5.3 Ablation Study
In this section, we conduct ablation studies for LANTERN. In Section 5.3.1, we assess the impact of the metric used to measure latent proximity on performance. Then, in Section 5.3.2, we provide an ablation study on the effect of the metric for measuring the distance from the modified probability distribution.
@@ 5.3.1 Nearest Latent Selection
Table 3: Ablation study for latent proximity measure and probability distribution distance metrics on LlamaGen-XL Stage I model under sampling ($\tau=1$). For latent proximity measures, $\ell_{2}$ distance, cosine similarity, and random selection are used, and TVD and JSD are used as probability distribution distance metrics.
@@
JSD ($\delta=0.13$) 2.09 17.48 0.3206
To explore the impact of various metrics for measuring latent proximity, we conduct an ablation study using representative distance metrics commonly used to measure latent proximity. Table 3 summarizes the comparisons among different strategies for selecting the nearest latent tokens, including $\ell_{2}$ distance, cosine similarity, and random selection under sampling ($\tau=1$). This experiment aims to assess the role of proximity-based selection in token aggregation, with $k=1000$ used across all methods.
As shown in Table 3, the random selection significantly underperforms in terms of acceleration, achieving only a $1.26$ mean accepted length, which is notably lower than both $\ell_{2}$ distance and cosine similarity. Additionally, the random selection shows inferior acceleration compared to the $\ell_{2}$ distance with $\delta=0.05$, while maintaining a similar FID, which highlights the importance of token selection based on latent proximity. Comparing $\ell_{2}$ distance and cosine similarity, both methods demonstrate nearly identical performance in terms of mean accepted length, FID, and CLIP score, suggesting robustness and flexibility in our approach to proximity measurement. These results emphasize that selecting tokens based on latent proximity is critical for achieving a balance between acceleration and image quality, offering a clear advantage over random selection. An analysis on the size of latent proximity set is presented in Appendix G.
@@ 5.3.2 Distance Between Probability Distribution
To ensure that the modified target distribution remains within an acceptable range of divergence from the original, we introduce $\delta$ as an upper bound for distributional divergence. To validate the impact of the divergence metric, we evaluate two different metrics to measure the divergence: Total Variation Distance (TVD) and Jensen-Shannon Divergence (JSD). Kullback-Leibler Divergence (KLD) is not used as it is asymmetric and not a valid metric for distance in a mathematical sense.
To compare the effectiveness of these distance metrics, we fix $k=1000$ and adjust $\delta$ to achieve similar mean accepted lengths across the different divergence metrics. The results shown in Table 3 demonstrate that, although the two metrics require different $\delta$ for the same level of acceleration (measured by mean accepted length), they produce nearly identical image quality metrics at comparable acceleration levels.
These results confirm that our method consistently functions as a robust trade-off controller regardless of the chosen distance metric. Since the difference in performance between two metrics is marginal, we opt to use TVD, as it is computationally lighter and thus more efficient for large-scale implementations. The detailed analysis on the latency differences between TVD and JSD is provided in Appendix F.3.
@@ 6 Conclusions
Acknowledgments
This work was partly supported by Institute for Information & communications Technology Promotion(IITP) (No.RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST), No.2022-0-00713, Meta-learning applicable to real-world problems, No.RS-2024-00457882, AI Research Hub Project) and National Research Foundation of Korea (NRF) (No.RS-2023- 00209060, A Study on Optimization and Network Interpretation Method for Large-Scale Machine Learning) grant funded by the Korea government (MSIT). This research was supported in part by the NAVER-Intel Co-Lab. The work was conducted by KAIST and reviewed by both NAVER and Intel.
@@ Appendix A Related Works
Visual auto-regressive models
The development of auto-regressive (AR) models in language modeling (Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023) has paved the way for AR-based approaches in generative vision tasks, where images are tokenized into discrete tokens on a 2D grid and processed as a unidirectional sequence. Early studies (Esser et al., 2021; Yu et al., 2021; Lee et al., 2022) explored different sequence orders such as row-major raster scans, spirals, and z-curves for image generation. However, these approaches often suffer from computational inefficiencies and underperform compared to diffusion models (Ho et al., 2020; Song et al., 2022; Rombach et al., 2022).
Building upon fundamental mechanism of visual AR models from DALL-E (Ramesh et al., 2021) and CogView (Ding et al., 2021), where images are discretized into tokens using vector-quantized VAEs (VQVAE) (Van Den Oord et al., 2017) or VQGANs (Esser et al., 2021) for next-token prediction, recently, visual AR models have emerged as strong competitors to diffusion models. For instance, Parti (Yu et al., 2022) and LlamaGen (Sun et al., 2024) utilize a verbose encoder-decoder architecture that incorporates frozen T5 (Raffel et al., 2020) text features via cross-attention or prefix-filling methods, drawing on insights from Imagen (Saharia et al., 2022). More recently, Chameleon (Team, 2024) has proposed a unified AR framework that enables fully token-based representations for both image and text modalities. Additionally, model like Anole (Chern et al., 2024), which build on Chameleon, further enhance image generation capabilities, demonstrating versatility across diverse tasks. In this work, we conduct experiments on state-of-the-art visual AR models—LlamaGen and Anole—for text-to-image generation to validate the effectiveness of our acceleration scheme.
Speculative decoding for AR models
The foundational concept of speculative decoding was first introduced by Leviathan et al. (2023) to address the sequential constraints of the AR framework in language modeling. Unlike the standard AR modeling, which generates one token per forward pass, speculative decoding aims to generate multiple tokens in a single forward pass by speculating a series of tokens, referred to as draft tokens. To generate draft tokens efficiently, most speculative decoding methods rely on a separate, typically smaller, model called the drafter model. Since generating draft tokens with the drafter model is much faster than the target model, this approach can reduce the overall token generation time.
In Leviathan et al. (2023), the drafter model is a smaller version within the same architecture family, trained on similar data and objectives as the target model. An alternative approach, DistillSpec (Zhou et al., ), uses knowledge distillation to develop efficient drafter models. Recognizing that the latency of generating draft tokens is crucial for accelerating, recent works have explored lightweight designs for drafter models. Medusa (Cai et al., 2024), for instance, employs multiple separate language model heads as drafter models instead of fully independent models and leverages tree drafting and decoding to enhance the chance of accepting draft tokens. Hydra (Ankner et al., 2024) builds on Medusa by incorporating sequential dependency, yielding additional acceleration benefits. EAGLE (Li et al., 2024b; a) extends these approaches by introducing feature uncertainty in drafting and utilizing a single decoder layer as drafter models to incorporate better sequential dependency. Furthermore, they have achieved state-of-the-art performance in speculative decoding for language modeling by dynamically refining the tree structure for tree drafting.
Appendix B Experimental Details
B.1 Experiments for Motivating Examples
For the experiment on naïve application shown in Figure 2(a), we utilize LlamaGen-L (Sun et al., 2024), an MLP with two linear layers, and a single decoder layer as the drafters for Speculative Decoding (Leviathan et al., 2023), Medusa (Cai et al., 2024), and EAGLE-2 (Li et al., 2024a), respectively. To facilitate Speculative Decoding, which requires a smaller-sized model with the same architecture as the main model and trained on the same dataset, we employ a class-conditioned LlamaGen-3B model instead of text-conditioned models. This is because text-conditioned models do not provide multiple model sizes. For each ImageNet class (Deng et al., 2009), we generate 100 images to measure the mean accepted length. Results on Vicuna-7B (Zheng et al., 2023) are taken directly from EAGLE-2.
The drafter test accuracies presented in Figure 2(b) correspond to the test accuracies of the drafters used in our main experiments (Table 2). Details on the training process for these drafters are provided in the following section. Additionally, the test accuracies of the learned drafter for Vicuna-13B are sourced from EAGLE-2.
For the experiments on average top-$1$ and top-$10$ probabilities shown in Figure 2(c), we generate 80 responses from Vicuna-7B on the MT-Bench dataset (Zheng et al., 2023) and 1000 images from LlamaGen-XL and Anole based on MS-COCO validation captions (Lin et al., 2014).
B.2 Experiments for Main Results
To train the text-conditional model’s drafter, we sampled 100k images in LAION-COCO dataset (Chuhmann et al., 2022), which is used to train Stage I target model. We used the same amount of image sampled in ImageNet (Deng et al., 2009) dataset to train the class-conditional model’s drafter. For Anole (Chern et al., 2024), we use 118K images sampled with target model by MS-COCO (Lin et al., 2014) train caption to train drafter. We used a single-layer decoder with the same structure as the target model in the same manner as EAGLE. During training, 5% of data is set to be held out validation dataset.
Since LlamaGen (Sun et al., 2024) and Anole (Chern et al., 2024) use classifier-free guidance (Ho & Salimans, 2021) to generate images, we trained our drafter to both learn conditioned input and null-conditioned input. To do so, we dropped 10% of conditional embedding during training, as same as target model training. The batch size is 16, and the base learning rate is $10^{-4}$. AdamW (Loshchilov & Hutter, 2019) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$ is used, and Linear learning rate scheduling with warm-up is used with 2000 warm-up steps. We select the best-performing model in terms of top-3 accuracy in the hold-out validation set for 20 epochs. In addition, Flan-T5 XL (Chung et al., 2022) is used to encode input text for text-conditional generation.
For LlamaGen (Sun et al., 2024) stage I and stage II, images are generated using a classifier-free guidance scale of 7.5 with top-p set to 1.0 and top-k set to 1000, which is the default generation configuration of LlamaGen official implementation for text-conditional image generation. For a class-conditional generation, the classifier-free guidance scale is set to 4.0, with the top-k sampling covering the entire vocabulary and the top-p sampling set to 1.0. For Anole (Chern et al., 2024), we use a classifier-free guidance scale of 3.0 with with top-k as 2000. For EAGLE-2 and our method, 60 candidate tokens are passed into the target model for each verification process.
Appendix C Detailed Analysis on Latent Proximity
C.1 Statistical Analysis on Latent Proximity
Table 4: Impact of replacing tokens with one of the $k$-th nearest tokens on image quality. FID and CLIP Score indicate degradation in image quality as $k$ increases.
Randomly Replaced by FID CLIP Score
one of $k$-th nearest token
Vanilla AR 25.06 0.3214
$k=50$ 26.88 0.3120
$k=100$ 30.76 0.3091
$k=1000$ 88.03 0.2715
Table 4 provides statistical evidence supporting our earlier qualitative observations (Figure 3) that token replacement does not lead to significant degradation in image quality, particularly for smaller values of $k$. This experiment was conducted using the LlamaGen-XL Stage I model (Sun et al., 2024) and MS-COCO 2017 validation captions (Lin et al., 2014), with FID and CLIP Score used as evaluation metrics.
As $k$ increases, the replaced token is chosen from a larger set of latent space neighbors, and its impact on image quality becomes more evident. For $k=50$, the FID increases slightly from 25.06 (Vanilla AR) to 26.88, and the CLIP Score decreases marginally from 0.3214 to 0.3120, indicating minimal degradation. Similarly, for $k=100$, the FID increases to 30.76, and the CLIP Score drops to 0.3091, showing that even with $k=100$, the image quality remains relatively stable and acceptable. However, for $k=1000$, a substantial decline is observed, with the FID increasing sharply to 88.03 and the CLIP Score dropping to 0.2715, underscoring the negative impact of replacing tokens with more distant neighbors.
These results corroborate our earlier qualitative findings, showing that token replacement up to $k=100$ maintains reasonable image quality, making it a viable approach in generative tasks. This demonstrates the robustness of the model under controlled token replacement scenarios.
C.2 Comparison of Latent Proximity in Visual and Language AR models
(a) A big teddy bear on a black motorcycle.
(b) A red and white train traveling along a mountain road.
Figure 6: Visualization of top-5 nearest token and top-5 farthest token from original token in LlamaGen-XL Stage II model (Sun et al., 2024).
Table 5: Top-5 nearest and farthest tokens from the original tokens based on $\ell_{2}$ distance of input embeddings in Vicuna-7B (Zheng et al., 2023). [OBJ] indicates the Unicode object replacement character.
Top-$k$ Token: hi Token: act
Nearest Farthest Nearest Farthest
[…]
2 _Mediabestanden [OBJ] oreferrer _Bruno
3 <0x4C> _infinitely <0x24> _Ernst
4 <0x6B> _firewall <0x71> _Santos
5 <0x49> _sooner <0x54> _firewall
To achieve a deeper understanding of latent proximity, we conducted additional analyses to examine the tendencies of nearest tokens in visual AR models and language models using LlamaGen-XL Stage II (Sun et al., 2024) and Vicuna-7B (Zheng et al., 2023). For the visual AR model, we visualized the five nearest tokens and the five farthest tokens based on their L2 distances in the latent space, as shown in Figure 6. However, for the language model, since each token corresponds to subwords, it is challenging to measure proximity directly using the token itself. Therefore, we conducted the same analysis using the $\ell_{2}$ distance of the input embeddings, with the results presented in Table 5.
This result indicates that in the visual AR model, tokens with high latent proximity are decoded into image patches with similar visual appearances, whereas tokens with low latent proximity correspond to image patches with distinctly different appearances. On the other hand, in the language model, the similarity based on the input embeddings did not reveal a clear relationship between the nearest and farthest tokens. These results support the conclusion that latent proximity serves as an effective metric for identifying visually similar tokens in visual AR models.
Appendix D Algorithms
D.1 Speculative Decoding with LANTERN